OpenStudy (anonymous):

A hamster ball has a radius of 21 centimeters. What is its volume written in terms of π? 12,348π cm3 6,946π cm3 588π cm3 9,261π cm3

OpenStudy (anonymous):

\[V_{Sphere} = \frac 43πr^3\]

OpenStudy (anonymous):

\[V_{Sphere} = (\frac 43)(π)(21^3)\]

OpenStudy (anonymous):

And here r is radius which is given as 21 cm Plug in the values of r = 21 and \(\pi = \frac{22}{7}\) and find the volume..

OpenStudy (anonymous):

Well, it says leave in terms of π, so you would just leave the π alone.

OpenStudy (anonymous):

Oh sorry I didn't see that..

OpenStudy (anonymous):

I got 9265.18 ?

OpenStudy (anonymous):

It is approx equal to D..

OpenStudy (anonymous):

Not quite...\[\frac{4 \times 9261 \times π}{3}\]

OpenStudy (anonymous):

Try dividing the 9261 and 3 first. Then multiply by 4.

OpenStudy (anonymous):

So the answer is A?

OpenStudy (anonymous):

Yup :)